OpenStudy (kmallette09):

The length of a rectangle is 2 feet longer than the width. The area of the rectangle is 8 square feet. Draw the graph of the function A(w) that represents this problem. Determine the maxima or minima.

OpenStudy (kmallette09):

will.h

OpenStudy (will.h):

no need, am here

OpenStudy (will.h):

i fanned you BTW

OpenStudy (kmallette09):

thanks

OpenStudy (will.h):

This question looks sick.

OpenStudy (kmallette09):

i know its tireing

OpenStudy (kmallette09):

tiring

OpenStudy (will.h):

what course is this. and what's the name of the lesson

OpenStudy (kmallette09):

appliations of quadratics

OpenStudy (will.h):

since length is 2ft longer than width then we know that L= w+2 since the are= 8 then we know that Length * width = 8 therefore W+2*W = 8 \[w^2 + 2w=8\] \[w^2+2w - 8 = 0\] Now we can factor it like this \[(w+4) (w-2) =0\] Now either \[W+4 = 0\] or \[W-2 = 0\] solve for the 1st one \[w+4 = 0\] \[w=-4\] which don't make sense because dimensions cannot be negative. solve the other one \[w-2 = 0\] \[w=2\] So the only reasonable width is 2 now substitute the width in the following equation to find the length. L=w+2 since we know that width = 2 then L= 2+2 L=4 Therefore the length is 4 and the width is 2.

OpenStudy (kmallette09):

thanks I would ask u for more help but u said u were tired

OpenStudy (will.h):

i can help with some more. maximum 2 more questions.

OpenStudy (kmallette09):

k